# Spontaneous Magnon Decays from Nonrelativistic Time-Reversal Symmetry Breaking in Altermagnets

### v1

@@
The isotropic Heisenberg model, with its spin-rotation symmetry, provides a reasonable approximation for quantum magnets when relativistic effects are negligible. In real materials, relativistic corrections such as magnetocrystalline anisotropy, $\lambda\sum_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}S_{{\boldsymbol{r}}}^{z}S_{{\boldsymbol{r}}^{\prime}}^{z}$, introduce perturbations that affect magnon stability. For easy-axis anisotropy ($\lambda<0$), relevant for altermagnetic candidates like MnF2, CoF2, and $\alpha$-Fe2O3 [19, 22], both Goldstone modes acquire a gap, growing as $\sqrt{\lambda J}$ due to the exchange enhancement. As $\lambda$ increases, the decay phase space shrinks, and beyond a critical threshold $\lambda^{\star}$, all magnons become stable, as expected in the Ising limit where the Néel ground state is an exact eigenstate. This provides a straightforward experimental criterion: for instance, neutron scattering data [31] set an upper bound of $\sim 120\,\mu$eV for the magnon splitting of MnF2, while its long-wavelength gap exceeds 111\,1meV, precluding three-magnon decays.
@@
Beyond global anisotropy, relativistic interactions such as Dzyaloshinskii-Moriya and pseudo-dipolar exchange can break spin-rotation symmetry, destabilizing magnetic order and enabling additional two-magnon decay processes [12, 13, 14, 15, 16, 17]. External magnetic fields can induce similar effects [10, 18]. These interactions contribute to magnon damping already at order $1/S$ [3], potentially overshadowing intrinsic $1/S^{2}$ many-body effects characteristic of altermagnets.
For an ideal platform to study pristine altermagnetic many-body magnon physics, the material should exhibit minimal relativistic corrections, or at most weak easy-axis anisotropy, to preserve approximate spin conservation. A quasi-two-dimensional structure and small spin quantum number $S$ is desirable to enhance quantum fluctuations, and a d-wave altermagnetic character leads to the most pronounced effects. The material should also be insulating to suppress electron-magnon damping, which also exhibits even-parity symmetries beyond s-wave [22, 33].
[…]
In discussions of altermagnets, their dual nature is often emphasized: they combine features of ferromagnets—such as spin splitting and time-reversal symmetry breaking—with those of antiferromagnets, including a compensated net moment [47]. The stability of the gapless, long-wavelength magnons discussed here reflects this hybrid character. Class-(i) magnets stabilize their gapless magnons via s-wave spin splitting, but their gapped branch is universally unstable. In contrast, class-(ii) magnets lack protective spin splitting, yet the curvature of their magnon dispersion can ensure stability under specific conditions. Class-(iii) magnets—altermagnets—present a combination of these mechanisms. At nodal planes in momentum space, where spin splitting vanishes, their magnons behave like those of class-(ii) magnets, albeit with corrections for d-wave altermagnets. However, in directions with spin-split magnon bands, the lower branch is stabilized in a manner reminiscent of class-(i) magnets for d-wave altermagnets, but not for g-wave and i-wave cases. As in class-(i) magnets, the upper branch remains universally unstable for that direction.
Beyond their significance in magnetism, altermagnets occupy a distinctive position in the broader landscape of spontaneous quasiparticle decay. Their continuous SO$(2)$ spin-space rotation symmetry makes spin-conserving three-particle decay the dominant instability, while two-particle decay is typically the main mechanism in other systems, such as phonons [5, 7] or magnons in relativistic and noncollinear magnets [3]. In collinear magnets, two-particle decay can also be suppressed by a discrete $C_{2}$ spin rotation symmetry [11, 3, 48], but in this case, three-particle decays are not constrained to conserve spin. Intriguingly, even continuous spin-space symmetries do not always suppress two-particle decay in all types of magnetic quasiparticles. For example, in quantum spin-gap magnets with singlet ground states, two-triplon decay remains allowed despite spin-rotation symmetry, as antisymmetric two-triplon combinations can still conserve total spin [49]. In these systems, additional real-space symmetries, such as reflections, may be necessary to suppress two-particle decays [50, 3]. Therefore, altermagnets present an intriguing platform for studying symmetry-constrained many-body dynamics.
Furthermore, while systems with multiple Goldstone modes—such as phonons in crystals or magnons in noncollinear magnets [3]—typically allow decay from faster to slower modes, this is generally driven by differences in their linear dispersion, for instance, by the velocity mismatch between transverse and longitudinal phonons. In altermagnets, however, the splitting between the modes is at least quadratic (d-wave), fundamentally altering the long-wavelength decay kinematics. Additionally, the even-parity nature of this splitting introduces a unique “island of stability” in magnon lifetimes. These findings highlight the exceptional many-body physics of altermagnets, signaling a marked departure from conventional paradigms in quantum magnetism and quasiparticle decay in condensed matter systems.
@@ Acknowledgements.
This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project No. 504261060 (Emmy Noether Programme) and within the Transregional Collaborative Research Center TRR 173/3 - 268565370 “Spin$+$X” (Project A03), by JSPS KAKENHI (Grants No. 20H00337, No. 22K14008, No. 23H04522, and No. 24H02231), JST CREST (Grant No. JPMJCR20T1), and by Waseda University Grant for Special Research Project (Grant No. 2024C-153). R. E. was supported by a Grant-in-Aid for JSPS Fellows (Grant No. 23KJ2047) and the JSPS Overseas Challenge Program for Young Researchers FY2024. M. G. acknowledges support by the Theory of Quantum Matter Unit of the Okinawa Institute of Science and Technology Graduate University (OIST), and by the Scientific Computing section of the Research Support Division at OIST for providing the HPC resources. A part of numerical simulations was carried out at the Supercomputer Center, Institute of Solid State Physics, University of Tokyo. The work of A. L. C. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award No. DE-SC0021221.
Methods
Spin-wave expansion for collinear isotropic magnets
[…]
@@ Methods
we map the spin operators onto sublattice bosons with annihilators $a_{{\boldsymbol{r}}}$ and $b_{{\boldsymbol{r}}}$ and creators $a^{\dagger}_{{\boldsymbol{r}}}$ and $b^{\dagger}_{{\boldsymbol{r}}}$. By expanding the square root in Eq. (14)—assuming ground state expectation values of $a^{\dagger}_{{\boldsymbol{r}}}a_{{\boldsymbol{r}}}$ and $b^{\dagger}_{{\boldsymbol{r}}}b_{{\boldsymbol{r}}}$ small compared to $2S_{\text{A/B}}$—we obtain an infinite expansion $H=H_{0}+H_{2}+H_{4}+H_{6}+\ldots$, where the subscript denotes the power in bosonic operators. There are no terms with an odd power of bosons because of the residual SO$(2)$ symmetry around the collinear axis. (More generally, a less stringent requirement for the absence of odd terms in the spin-wave expansion of collinear magnets is a $C_{2}$ symmetry around the collinear axis [11, 3, 48]. In this case, however, the even terms do not have to conserve spin.) $H_{0}$ is the classical ground state energy that is of no further interest. $H_{2}$ encodes the harmonic theory of noninteracting magnons, and $H_{4}$ and beyond captures magnon-magnon interactions. In momentum space, the harmonic theory is block diagonal,
@@
where the two blocks correspond to opposite spin quantum numbers ±$\pm 1$. Here, $\delta H_{0}$ is a number providing a correction to the ground state energy, and the two-by-two Hamilton matrix blocks read
@@
[…]
@@
[…]
where we have suppressed momentum labels. The full expression in given in the Supplementary Note 2. Notably, none of the terms in Eq. (18) change the total spin—the total spin of the destroyed bosons equals that of those created. While Ref. [23] suggests the presence of terms like $\alpha^{\dagger}\alpha^{\dagger}\alpha^{\dagger}\beta$, such combinations do not actually appear. The same observation holds for $H_{6}$ and higher-order terms and is a consequence of spin conservation as discussed above.
@@
The interaction-corrected magnon dispersion $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})$ is obtained from the poles of the retarded Green’s function, $G_{\pm}({\boldsymbol{k}},\omega)=[\omega+\mathrm{i}0^{+}-\epsilon_{\pm}({\boldsymbol{k}})-\Sigma^{(1/S)}_{\pm}({\boldsymbol{k}},\omega)-\Sigma^{(1/S^{2})}_{\pm}({\boldsymbol{k}},\omega)]^{-1}$, where the relevant zero-temperature single-magnon self-energies, $\Sigma^{(1/S)}_{\pm}({\boldsymbol{k}},\omega)$ and $\Sigma^{(1/S^{2})}_{\pm}({\boldsymbol{k}},\omega)$, are given in the Supplementary Note 2. Within an on-shell calculation, the corrected spectrum is found to read
@@
For the off-shell calculation presented in Fig. 5b, we have plotted the single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)=-\text{Im}[G_{+}({\boldsymbol{k}},\omega)+G_{-}({\boldsymbol{k}},\omega)]/\pi$. The resulting spectrum is no longer consistent within $1/S$ and captures some nonperturbative effects, such as non-Lorenzian lineshapes. This inconcsistency in $1/S$ generates a problem for the Goldstone mode that gets pushed to negative energies. Numerically, we found it necessary to suppress one particular family of self-energy diagrams to avoid this problem; this neglection has almost no visible effects on the spectrum away from the Brillouin zone origin as the magnitude of these self-energies is tiny (see Supplementary Note 2 for further details).
Linear spin-waves in checkerboard d-wave altermagnet
@@
Matrix product operator based time evolution
@@
Figure 6: Cylinder geometries and accessible momenta in reciprocal space in nonperturbative DMRG$+$tMPO simulations. (a) Regular square lattice with a circumference of $L_{y}=4$ unit cells. (b) $\pi/4$-rotated lattice with $L_{y}=3$. (c) Brillouin zone and accessible momenta cuts of regular (solid) and rotated lattice (dashed). The red line illustrates the Brillouin zone path used in Fig. 5a. The segment from $(-\frac{\pi}{2},\frac{\pi}{2})$ to $(-\pi,\pi)$ is taken from the path $(\frac{\pi}{2},\frac{\pi}{2})$ to $(\pi,\pi)$. Although these paths are not related by symmetry and have opposite altermagnetic splitting, the transverse dynamical structure factor $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ is identical.
@@
where $\gamma=\{x,y,z\}$ and $C^{\gamma\gamma}_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}(t)=\langle S^{\gamma}_{{\boldsymbol{r}}}(t)S^{\gamma}_{{\boldsymbol{r}}^{\prime}}(0)\rangle$ are the dynamical spin-spin correlations. The protocol is as follows: (i) $C^{\gamma\gamma}_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}(t)$ is Fourier transformed in space providing $C^{\gamma\gamma}({\boldsymbol{k}},t)$, (ii) $C^{\gamma\gamma}({\boldsymbol{k}},t)$ gets extended in time using linear predictive coding [61, 62, 63, 64], and convoluted with a Gaussian of width $\sigma_{t}\approx 11.16$ to suppress ringing from the box function of the finite time window, and (iii) in doing a Fourier transform in time and coherent summations over the two sublattices of the magnetic unit cell, we obtain $\mathcal{S}_{\gamma\gamma}({\boldsymbol{k}},\omega)$ as in Eq. (28). The convolution with a Gaussian in (ii) results in a Gaussian broadening in $\mathcal{S}_{\gamma\gamma}(\bm{k},\omega)$ of width $\sigma_{\omega}\approx 0.0896$. The final $\mathcal{S}_{\perp}(\bm{k},\omega)=\mathcal{S}_{xx}(\bm{k},\omega)+\mathcal{S}_{yy}(\bm{k},\omega)$ is plotted in Fig. 5a.
@@
Our simulations complement those of Ref. [23]. Here, we remark on the difference to the approach taken in Ref. [23] and its implications. We compute the ground state on an infinite cylinder and perform the time-evolution on a finite cylinder segment embedded in an infinite cylinder. This is to be contrasted to the finite length, $L_{x}=80$, cylinder in Ref. [23]. Therefore, we are less affected by finite-size effects. In turn, we add a finite XXZ anisotropy, $\lambda=0.02$, to control magnetization, convergence and computational resources, while retaining interpretability of the gap. For comparison, Ref. [23] has a spin gap of $\varepsilon\approx 0.5$, while at $\lambda=0.02$ we obtain $\varepsilon\approx 0.23$. Furthermore, the ground state of the isotropic, $\lambda=0$, Heisenberg model on a finite-width cylinder does not break $\mathrm{SO}(3)$ spontaneously and develops a paramagnetic, singlet-like ground state instead. As a result, $\mathcal{S}_{\mu\mu}({\boldsymbol{k}},\omega)$ with $\mu=x,y,z$ is isotropic and picks up single-magnon and two-magnon contributions. In our case, an XXZ anisotropy of $\lambda=0.02$ ensures finite sublattice magnetization, which enables us to separate the transverse $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ from the longitudinal $\mathcal{S}_{zz}({\boldsymbol{k}},\omega)$, the latter of which we are presenting in Supplementary Note 5. However, as $\mathcal{S}_{zz}({\boldsymbol{k}},\omega)$ is proportional to the two-magnon sector rather than single particles and therefore does not directly reflect quasiparticle damping, we propose measurement of $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ in experiments.
Magnon damping estimates for $\kappa$-Cl-like altermagnets
@@
Figure 7: Spontaneous magnon decay in $\kappa$-Cl-like organic altermagnets. (a,b) Effective spin model of $\kappa$-Cl on the square lattice derived in Ref. [21]. Magenta and cyan arrows with ellipsoids respectively denote the collinear ground state $(\pi,\pi)$ Néel order. Directions of the ellipsoids correspond to the molecular orientations, which is the origin of altermagnetism in this material. White and green bonds in (a) indicate the nearest-neighbor and the next-nearest-neighbor exchange interactions $J$ and $J^{\prime}$, respectively. In (b), due to the differing molecular orientations between the two sublattices, the third-nearest neighbor exchange interactions along the bonds parallel and perpendicular to the orientation direction are non-equivalent. Here, only the latter is denoted as $K$ and the former is neglected. (c) Renormalized magnon spectrum (lines with dot markers) compared to the linear magnon spectrum (solid lines) for $J=80\,$meV, $J^{\prime}=20\,$meV, $S=1/2$ and $K=20\,$meV. The value of $K$ sets the magnitude of the altermagnetic splitting and is approximately ten times larger than what is predicted for $\kappa$-Cl [21]. (d) Magnon damping $\Gamma_{\pm}({\boldsymbol{k}})$ along momentum directions with maximal band splitting.
@@
Here, the local spin operators ${\boldsymbol{S}}_{{\boldsymbol{r}}}$ of length $S=1/2$ are associated with molecular dimers, which are indicated by ellipsoids in Fig. 7a. $J$ and $J^{\prime}$ are the first and second-nearest-neighbor exchange interaction, respectively. $K$ is a third-nearest neighbor interaction along those third-neighbor bonds that align with the short axis of the dimers, see Fig. 7b. This third-nearest neighbor coupling causes the altermagnetic splitting in contrast to the model in Eq. (8), where it derives from second-neighbor coupling. As a result, the magnon splitting is maximal at the midpoint of the line from $(0,0)$ to $(\pi,0)$ or to $(0,\pi)$, see Fig. 7c, but it is zero at the entire boundary of the Brillouin zone. By carrying out a nonlinear spin-wave analysis identical to the one outlined above, we find that for the parameters predicted in Ref. [21], that is, $J=80\,$meV, $J^{\prime}=20\,$meV, and $K=2\,$meV, the maximal magnon damping is smaller than $1\,\mu$eV. However, if $K$ is increased by a factor of ten, we obtain a maximal damping of $\Gamma\sim 6\,$meV, as shown in Fig. 7d.
@@
±$\quad\pm\Delta\cos(2\varphi_{\boldsymbol{k}})|{\boldsymbol{k}}|^{2}.$ (30)
@@
We find that $|\nu_{2\varphi}|\ll\nu$ and $|\kappa_{2\varphi}|$, $|\kappa_{4\varphi}|\ll|\kappa|$ are satisfied. The angular dependence of the velocity $\nu_{2\varphi}$ and the cubic nonlinearity $\kappa_{2\varphi}$ stems from the next-nearest-neighbor exchange interaction $J^{\prime}$, which breaks the four-fold rotation symmetry of the square lattice. Since the crystal structure of $\kappa$-Cl belongs to the crystallographic Laue group $mmm$, which does not preserve four-fold rotational symmetry, both $\nu_{2\varphi}$ and $\kappa_{2\varphi}$ originate purely from lattice-geometric factors. We emphasize that a finite $\nu_{2\varphi}$ does not disrupt the scaling theory of $\Gamma$ in the long-wavelength limit.

### v2

@@ Discussion
The isotropic Heisenberg model describes quantum magnets when relativistic effects are negligible. In real materials, relativistic corrections such as magnetocrystalline anisotropy, $\lambda\sum_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}S_{{\boldsymbol{r}}}^{z}S_{{\boldsymbol{r}}^{\prime}}^{z}$, influence magnon stability. For easy-axis anisotropy ($\lambda<0$), relevant for altermagnetic candidates like MnF2, CoF2, and $\alpha$-Fe2O3 [19, 22, 33], both Goldstone modes acquire a gap scaling as $\sqrt{\lambda J}$. Beyond a critical anisotropy $\lambda^{\star}$, magnons become fully stable, a situation likely realized in MnF2 [34]. Experimental evidence for altermagnetic magnon spin splitting exists for MnTe [35], which shows a $2\,\text{meV}$ g-wave splitting over a $36\,\text{meV}$ bandwidth. However, its easy-plane anisotropy gaps one Goldstone mode by $4\,\text{meV}$ and violates spin conservation, requiring separate analysis. So do relativistic interactions such as Dzyaloshinskii-Moriya and pseudo-dipolar exchange (and field-induced canting) that enable order-$1/S$ two-magnon decay processes [10, 3, 12, 13, 18, 14, 15, 16, 17], which potentially overshadow intrinsic $1/S^{2}$ many-body effects characteristic of altermagnets.
@@
Altermagnets occupy a unique position in the broader context of spontaneous quasiparticle decay. Their continuous SO(2)(2) symmetry makes spin-conserving three-particle decay the dominant instability, unlike other systems where two-particle decay prevails, such as phonons [5, 7] or magnons in relativistic and noncollinear magnets [3]. Note the difference to collinear magnets with a discrete $C_{2}$ spin symmetry whose two-particle decay is also suppressed but their three-particle decays need not conserve spin [11, 3, 50]. Notably, continuous spin symmetry does not always prevent two-particle decay—for instance, in quantum spin-gap magnets, two-triplon decay is allowed despite spin conservation [51], requiring additional real-space symmetries for suppression [52, 3]. This makes altermagnets a compelling platform to explore symmetry-constrained many-body dynamics.
In systems with multiple Goldstone modes, such as phonons [3], decay from faster to slower modes usually depends on linear dispersion differences. In altermagnets, however, the mode splitting is at least quadratic, fundamentally altering decay kinematics and giving rise to a unique “island of stability” in magnon lifetimes. These features underscore the exceptional many-body physics of altermagnets, marking a significant departure from conventional paradigms in quantum magnetism and quasiparticle decay.
@@
Data availability
All data and codes are available from the authors upon reasonable request.
@@ Acknowledgements.
This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project No. 504261060 (Emmy Noether Programme) and within the Transregional Collaborative Research Center TRR 173/3 - 268565370 “Spin++X” (Project A03), by JSPS KAKENHI (Grants No. 22K14008, No. 24H02231, and No. 25H00611), JST CREST (Grant No. JPMJCR20T1), and by Waseda University Grant for Special Research Project (Grant No. 2025C-133). R. E. was supported by a Grant-in-Aid for JSPS Fellows (Grant No. 23KJ2047) and the JSPS Overseas Challenge Program for Young Researchers FY2024. M. G. acknowledges support by the Theory of Quantum Matter Unit of the Okinawa Institute of Science and Technology Graduate University (OIST). The work of A. L. C. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award No. DE-SC0021221. A part of numerical simulations was carried out at the Supercomputer Center, Institute of Solid State Physics, University of Tokyo. M. G. acknowledges support by the Scientific Computing section of the Research Support Division at OIST for providing the HPC resources.
Appendix A: Spin-wave expansion for collinear isotropic magnets
We assume that the classical ground state $|\text{CGS}\rangle$ of the nonrelativistic SO(3)(3)-symmetric Heisenberg model in Eq. (1) exhibits spontaneous symmetry breaking with collinear order, such that the magnet falls in one of the three classes mentioned in the main text. Without loss of generality, we fix the collinear order pointing along the $z$-direction, such that $|\text{CGS}\rangle$ becomes a product state of fully (positively or negatively) polarized eigenstates of the local spin-$z$ operator $S^{z}_{{\boldsymbol{r}}}$, i.e., the uniform state for ferromagnets and the Néel state for antiferromagnets and altermagnets. There is still a residual SO(2)(2) symmetry around the collinear axis, suggesting that the total spin-$z$ operator $S^{z}=\sum_{{\boldsymbol{r}}}S_{{\boldsymbol{r}}}^{z}$ commutes with $H$. However, $|\text{CGS}\rangle$ is only an eigenstate of $S^{z}$ but not of $H$. The situation is particularly pronounced for compensated magnets because their exact ground state is a fully isotropic spin singlet state [54]. However, as argued by Anderson [54], the exact ground state is irrelevant because one can build a state with finite sublattice magnetizations as a superposition of states that are degenerate with the ground state in the thermodynamic limit. The point of a spin-wave expansion is then to approximate such a state perturbatively, starting from $|\text{CGS}\rangle$. Since $|\text{CGS}\rangle$ is an eigenstate of $S^{z}$, so are all excited states encountered in the spin-wave expansion, which are built upon $|\text{CGS}\rangle$ by acting on it with $S_{{\boldsymbol{r}}}^{\pm}=S_{{\boldsymbol{r}}}^{x}\pm\mathrm{i}S_{{\boldsymbol{r}}}^{y}$. As a particular result, magnons—obtained by a single action of $S_{{\boldsymbol{r}}}^{\pm}$—carry a definite spin quantum number of $\pm 1$ ($\hbar=1$) relative to the ground state, and the interactions among them may not change the spin.
@@
we map the spin operators onto sublattice bosons with annihilators $a_{{\boldsymbol{r}}}$ and $b_{{\boldsymbol{r}}}$ and creators $a^{\dagger}_{{\boldsymbol{r}}}$ and $b^{\dagger}_{{\boldsymbol{r}}}$. By expanding the square root in Eq. (14)—assuming ground state expectation values of $a^{\dagger}_{{\boldsymbol{r}}}a_{{\boldsymbol{r}}}$ and $b^{\dagger}_{{\boldsymbol{r}}}b_{{\boldsymbol{r}}}$ small compared to $2S_{\text{A/B}}$—we obtain an infinite expansion $H=H_{0}+H_{2}+H_{4}+H_{6}+\ldots$, where the subscript denotes the power in bosonic operators. There are no terms with an odd power of bosons because of the residual SO(2)(2) symmetry around the collinear axis. (More generally, a less stringent requirement for the absence of odd terms in the spin-wave expansion of collinear magnets is a $C_{2}$ symmetry around the collinear axis [11, 3, 50]. In this case, however, the even terms do not have to conserve spin.) $H_{0}$ is the classical ground state energy that is of no further interest. $H_{2}$ encodes the harmonic theory of noninteracting magnons, and $H_{4}$ and beyond captures magnon-magnon interactions. In momentum space, the harmonic theory is block diagonal,
@@
where the two blocks correspond to opposite spin quantum numbers $\pm 1$. Here, $\delta H_{0}$ is a number providing a correction to the ground state energy, and the two-by-two Hamilton matrix blocks read
@@
The Bogoliubov vacuum $|\varnothing\rangle$, defined by $\alpha_{{\boldsymbol{k}}}|\varnothing\rangle=\beta_{{\boldsymbol{k}}}|\varnothing\rangle=0$, is the leading-order quantum correction to the classical ground state $|\text{CGS}\rangle$. It can be considered a two-mode squeezed vacuum of those sublattice-boson Fock states that carry the same number of bosons on sublattices A and B [55]. $|\varnothing\rangle$ is an eigenstate of $S^{z}$ with the same eigenvalue as $|\text{CGS}\rangle$. Likewise, the single-magnon states $\alpha^{\dagger}_{{\boldsymbol{k}}}|\varnothing\rangle$ and $\beta^{\dagger}_{{\boldsymbol{k}}}|\varnothing\rangle$ are eigenstates of $S^{z}$ because they are two-mode squeezed states of those sublattice-boson Fock states that carry $n+1$ bosons on sublattice A and $n$ on sublattice B (or vice versa) [55]. They carry spin relative to the vacuum identical to that of the sublattice-boson Fock states, that is (n+1)n=+1(n+1)-n=+1 or n(n+1)=1n-(n+1)=-1. The spin of magnons in collinear magnets is alternatively referred to as their handedness or chirality [22], or polarization [56].
@@
$\begin{aligned} H_{4}&=W^{(1)}\alpha^{\dagger}\alpha^{\dagger}\alpha\alpha+W^{(2)}\beta^{\dagger}\beta^{\dagger}\beta\beta+W^{(3)}\alpha^{\dagger}\beta^{\dagger}\alpha\beta\\ &\quad+W^{(4)}\alpha^{\dagger}\alpha^{\dagger}\beta^{\dagger}\alpha+W^{(5)}\alpha^{\dagger}\beta^{\dagger}\beta^{\dagger}\beta+W^{(6)}\alpha^{\dagger}\alpha^{\dagger}\beta^{\dagger}\beta^{\dagger}\\ &\quad+W^{(7)}\alpha^{\dagger}\alpha\alpha\beta+W^{(8)}\beta^{\dagger}\alpha\beta\beta+W^{(9)}\alpha\alpha\beta\beta,\end{aligned}$ (18)
where we have suppressed momentum labels. The full expression in given in Supplementary Material II. Notably, none of the terms in Eq. (18) change the total spin—the total spin of the destroyed bosons equals that of those created. While Ref. [25] suggests the presence of terms like $\alpha^{\dagger}\alpha^{\dagger}\alpha^{\dagger}\beta$, such combinations do not actually appear. The same observation holds for $H_{6}$ and higher-order terms and is a consequence of spin conservation as discussed above.
@@
The interaction-corrected magnon dispersion $\tilde{\epsilon}_{\pm}({\boldsymbol{k}})$ is obtained from the poles of the retarded Green’s function, $G_{\pm}({\boldsymbol{k}},\omega)=[\omega+\mathrm{i}0^{+}-\epsilon_{\pm}({\boldsymbol{k}})-\Sigma^{(1/S)}_{\pm}({\boldsymbol{k}},\omega)-\Sigma^{(1/S^{2})}_{\pm}({\boldsymbol{k}},\omega)]^{-1}$, where the relevant zero-temperature single-magnon self-energies, $\Sigma^{(1/S)}_{\pm}({\boldsymbol{k}},\omega)$ and $\Sigma^{(1/S^{2})}_{\pm}({\boldsymbol{k}},\omega)$, are given in Supplementary Material II. Within an on-shell calculation, the corrected spectrum is found to read
@@
For the off-shell calculation presented in Fig. 5b, we have plotted the single-particle spectral function $\mathcal{A}({\boldsymbol{k}},\omega)=-\text{Im}[G_{+}({\boldsymbol{k}},\omega)+G_{-}({\boldsymbol{k}},\omega)]/\pi$. The resulting spectrum is no longer consistent within $1/S$ and captures some nonperturbative effects, such as non-Lorenzian lineshapes. This inconcsistency in $1/S$ generates a problem for the Goldstone mode that gets pushed to negative energies. Numerically, we found it necessary to suppress one particular family of self-energy diagrams to avoid this problem; this neglection has almost no visible effects on the spectrum away from the Brillouin zone origin as the magnitude of these self-energies is tiny (see Supplementary Material II for further details).
Appendix B: Linear spin-waves in checkerboard d-wave altermagnet
@@
Appendix C: Matrix product operator based time evolution
@@
Figure 6: Cylinder geometries and accessible momenta in reciprocal space in nonperturbative DMRG++tMPO simulations. (a) Regular square lattice with a circumference of $L_{y}=4$ unit cells. (b) $\pi/4$-rotated lattice with $L_{y}=3$. (c) Brillouin zone and accessible momenta cuts of regular (solid) and rotated lattice (dashed). The red line illustrates the Brillouin zone path used in Fig. 5a. The segment from $(-\frac{\pi}{2},\frac{\pi}{2})$ to $(-\pi,\pi)$ is taken from the path $(\frac{\pi}{2},\frac{\pi}{2})$ to $(\pi,\pi)$. Although these paths are not related by symmetry and have opposite altermagnetic splitting, the transverse dynamical structure factor $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ is identical.
@@
where $\gamma=\{x,y,z\}$ and $C^{\gamma\gamma}_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}(t)=\langle S^{\gamma}_{{\boldsymbol{r}}}(t)S^{\gamma}_{{\boldsymbol{r}}^{\prime}}(0)\rangle$ are the dynamical spin-spin correlations. The protocol is as follows: (i) $C^{\gamma\gamma}_{{\boldsymbol{r}},{\boldsymbol{r}}^{\prime}}(t)$ is Fourier transformed in space providing $C^{\gamma\gamma}({\boldsymbol{k}},t)$, (ii) $C^{\gamma\gamma}({\boldsymbol{k}},t)$ gets extended in time using dynamical mode decomposition (DMD) [63, 64] and convoluted with a Gaussian of width $\sigma_{t}\approx 13.02$ to suppress ringing from the box function of the finite time window, and (iii) in doing a Fourier transform in time and coherent summations over the two sublattices of the magnetic unit cell, we obtain $\mathcal{S}_{\gamma\gamma}({\boldsymbol{k}},\omega)$ as in Eq. (28). The convolution with a Gaussian in (ii) results in a Gaussian broadening in $\mathcal{S}_{\gamma\gamma}(\bm{k},\omega)$ of width $\sigma_{\omega}\approx 0.0768$. The final $\mathcal{S}_{\perp}(\bm{k},\omega)=\mathcal{S}_{xx}(\bm{k},\omega)+\mathcal{S}_{yy}(\bm{k},\omega)$ is plotted in Fig. 5a.
@@
Our simulations complement those of Ref. [25]. Here, we remark on the difference to the approach taken in Ref. [25] and its implications. We compute the ground state on an infinite cylinder and perform the time-evolution on a finite cylinder segment embedded in an infinite cylinder. This is to be contrasted to the finite length, $L_{x}=80$, cylinder in Ref. [25]. Therefore, we are less affected by finite-size effects. In turn, we add a finite XXZ anisotropy, $\lambda=0.02$, to control magnetization, convergence and computational resources, while retaining interpretability of the gap. For comparison, Ref. [25] has a spin gap of $\varepsilon\approx 0.5$, while at $\lambda=0.02$ we obtain $\varepsilon\approx 0.23$. Furthermore, the ground state of the isotropic, $\lambda=0$, Heisenberg model on a finite-width cylinder does not break $\mathrm{SO}(3)$ spontaneously and develops a paramagnetic, singlet-like ground state instead. As a result, $\mathcal{S}_{\mu\mu}({\boldsymbol{k}},\omega)$ with $\mu=x,y,z$ is isotropic and picks up single-magnon and two-magnon contributions. In our case, an XXZ anisotropy of $\lambda=0.02$ ensures finite sublattice magnetization, which enables us to separate the transverse $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ from the longitudinal $\mathcal{S}_{zz}({\boldsymbol{k}},\omega)$, the latter of which we are presenting in Supplementary Material V. However, as $\mathcal{S}_{zz}({\boldsymbol{k}},\omega)$ is proportional to the two-magnon sector rather than single particles and therefore does not directly reflect quasiparticle damping, we propose measurement of $\mathcal{S}_{\perp}({\boldsymbol{k}},\omega)$ in experiments.
Appendix D: Magnon damping estimates for $\kappa$-Cl-like altermagnets
@@
Figure 7: Spontaneous magnon decay in $\kappa$-Cl-like organic altermagnets. (a,b) Effective spin model of $\kappa$-Cl on the square lattice derived in Ref. [21]. Magenta and cyan arrows with ellipsoids respectively denote the collinear ground state $(\pi,\pi)$ Néel order. Directions of the ellipsoids correspond to the molecular orientations, which is the origin of altermagnetism in this material. White and green bonds in (a) indicate the nearest-neighbor and the next-nearest-neighbor exchange interactions $J$ and $J^{\prime}$, respectively. In (b), due to the differing molecular orientations between the two sublattices, the third-nearest neighbor exchange interactions along the bonds parallel and perpendicular to the orientation direction are non-equivalent. Here, only the latter is denoted as $K$ and the former is neglected. (c) Renormalized magnon spectrum (lines with dot markers) compared to the linear magnon spectrum (solid lines) for J=80J=80\,meV, $J^{\prime}=20\,$meV, $S=1/2$ and K=20K=20\,meV. The value of $K$ sets the magnitude of the altermagnetic splitting and is approximately ten times larger than what is predicted for $\kappa$-Cl [21]. (d) Magnon damping $\Gamma_{\pm}({\boldsymbol{k}})$ along momentum directions with maximal band splitting.
@@
In our initial long-wavelength expansion of the magnon dispersion of altermagnets in Eq. (3) we have neglected the crystallographic influence on the velocity $v$ and the cubic nonlinearity $\kappa$. Taking them into account for $\kappa\textrm{-Cl}$, we obtain
@@
$\quad\pm\Delta\cos(2\varphi_{\boldsymbol{k}})|{\boldsymbol{k}}|^{2}.$ (30)
@@
We find that $|\nu_{2\varphi}|\ll\nu$ and $|\kappa_{2\varphi}|$, $|\kappa_{4\varphi}|\ll|\kappa|$ are satisfied. The angular dependence of the velocity $\nu_{2\varphi}$ and the cubic nonlinearity $\kappa_{2\varphi}$ stems from the next-nearest-neighbor exchange interaction $J^{\prime}$, which breaks the four-fold rotation symmetry of the square lattice. Since the crystal structure of $\kappa$-Cl belongs to the crystallographic Laue group mmmmmm, which does not preserve four-fold rotational symmetry, both $\nu_{2\varphi}$ and $\kappa_{2\varphi}$ originate purely from lattice-geometric factors. We emphasize that a finite $\nu_{2\varphi}$ does not disrupt the scaling theory of $\Gamma$ in the long-wavelength limit.